# Multi-Frequency Progressive Refinement for Learned Inverse Scattering

### v1

@@ Acknowledgements
OM, OT, VC, and RW gratefully acknowledge the support of AFOSR FA9550-18-1-0166 and NSF DMS-2023109. RW and YK gratefully acknowledge the support of DOE DE-SC0022232. The team gratefully acknowledges the support of the Margot and Tom Pritzker Foundation.
@@ Appendix A Distribution of Scattering Potentials
The random low-frequency backgrounds were generated by drawing random Fourier coefficients and filtering out the high frequencies using $\mathsf{LPF}_{4.0}$. The resulting background was transformed to Cartesian coordinates, and then shifted and scaled so the maximum value was $2.0$ and the minimum value was $0.0$. The background was truncated to $0.0$ outside of the disk of radius $0.4$. Three shapes were randomly chosen among equilaterial triangles, squares, and ellipses. The three shapes had randomly-chosen centers and rotations, constrained to be non-overlapping and fit inside the disk of radius $0.4$. The side lengths of the squares and triangles were uniformly sampled from $[0.1,0.15]$. The major axis lengths of the ellipses were uniformly sampled from $[0.1,0.15]$, and the minor axis lengths were uniformly sampled from $[0.05,0.1]$. Finally, $\mathsf{LPF}_{32\pi}$ was applied to the scattering potential.
@@ Appendix B Hyperparameter Search
B.1 MFISNet Models
We train all models using the Adam algorithm, with a batch size of $16$ samples. All of the FYNet, MFISNet-Parallel, MFISNet-Fused, and MFISNet-Refinement models tested have 3 1D convolutional layers followed by 3 2D convolutional layers; ReLU activations are used between layers. In the FYNet, MFISNet-Parallel and MFISNet-Refinement models, we use 1D and 2D convolutional kernels with $24$ channels following [Fan and Ying, 2022]; in the MFISNet-Fused models, we use 1D and 2D convolutional kernels with $24N_{k}$ channels to adjust for the increased input size. To train MFISNet-Fused, MFISNet-Parallel and MFISNet-Refinement, we search over a grid of architecture and optimization hyperparameters. We report the optimal hyperparameters we found in Tables 4, 7, 6, 5 and 8.
@@ B.1 MFISNet Models
1d kernel size 60 40
2d kernel size 7 5
Weight decay $1\times 10^{-3}$ $0.0$
@@
1d kernel size $20$ $20$ $20$ $40$ $20$ $20$ $20$ $20$
2d kernel size $5$ $5$ $7$ $5$ $5$ $5$ $7$ $7$
Weight decay $1\times 10^{-3}$ $0.0$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $0.0$ $1\times 10^{-3}$ $0.0$
Learning rate $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
LR decrease $0.25$ $0.25$ $0.25$ $0.25$ $0.25$ $0.25$ $0.25$ $0.25$
@@
1d kernel size $40$ $20$ $20$ $20$ $40$ $40$ $20$ $40$
2d kernel size $7$ $7$ $7$ $7$ $7$ $5$ $7$ $7$
Weight decay $1\times 10^{-3}$ $0.0$ $0.0$ $1\times 10^{-3}$ $1\times 10^{-3}$ $0.0$ $1\times 10^{-3}$ $0.0$
Learning rate $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$
@@
1d kernel size 40 20 20 60 40 20 20 60
2d kernel size 7 5 5 5 5 5 5 5
Weight decay $0.0$ $1\times 10^{-3}$ $1\times 10^{-3}$ $0.0$ $0.0$ $0.0$ $0.0$ $1\times 10^{-3}$
Learning Rate $1\times 10^{-3}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $5\times 10^{-4}$
@@
Hyperparameter Training Method
No Iterative Refinement No Homotopy through Frequency Our Method
1d kernel size $20$ $40$ $40$
2d kernel size $7$ $7$ $5$
Weight decay $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Learning rate $5\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-3}$
LR decrease N/A N/A $0.25$
$\mathbf{\gamma}$ N/A $1.1$ N/A
Table 8: Optimal hyperparameters for the MFISNet-Refinement models in the training method experiment (Section 5.5). Here, $\mathbf{\gamma}$ is the factor which weights different loss terms in the “No Homotopy through Frequency” condition (cf. (16)). We searched over values $\gamma=\{0.9,1.0,1.1\}$. All of these models were trained with incident frequencies $[2\pi,4\pi,8\pi,16\pi,32\pi]$ and $n=2000$ noiseless training samples.
@@ B.2 Wide-Band Butterfly Network
Appendix C Additional empirical results
In this section, we illustrate the predictions generated by the different models used in our experiments on four randomly-selected test samples from our dataset. For each prediction, we include the associated error plot. The visualizations are provided in Figure 8.
[…]
[width=]svg-inkscape/preds_compared_along_n_omega_jcp_appendix_0_svg-tex.pdf_tex
Figure 8: Sample predictions from four randomly-selected test samples.

### v2

@@ Acknowledgements
The authors would like to thank Borong Zhang for sharing his implementation of the Wide-Band Butterfly Network. OM, OT, VC, and RW gratefully acknowledge the support of AFOSR FA9550-18-1-0166 and NSF DMS-2023109. RW and YK gratefully acknowledge the support of DOE DE-SC0022232. YK gratefully acknowledges the support of NSF DMS-2339439. The team gratefully acknowledges the support of the Margot and Tom Pritzker Foundation.
@@ Appendix A Distribution of Scattering Potentials
The random low-frequency backgrounds were generated by drawing random Fourier coefficients and filtering out the high frequencies using $\mathsf{LPF}_{7.\bar{1}\pi}$. The resulting background was transformed to Cartesian coordinates, and then shifted and scaled so the maximum value was $2.0$ and the minimum value was $0.0$. The background was truncated to $0.0$ outside of the disk of radius $0.4$. Three shapes were randomly chosen among equilaterial triangles, squares, and ellipses. The three shapes had randomly-chosen centers and rotations, constrained to be non-overlapping and fit inside the disk of radius $0.4$. The side lengths of the squares and triangles were uniformly sampled from $[0.1,0.15]$. The major axis lengths of the ellipses were uniformly sampled from $[0.1,0.15]$, and the minor axis lengths were uniformly sampled from $[0.05,0.1]$. Finally, $\mathsf{LPF}_{32\pi}$ was applied to the scattering potential.
@@ Appendix B Hyperparameter Search
B.1 FYNet and MFISNet Models
We train all models using the Adam algorithm, with a batch size of $16$ samples. All of the FYNet, MFISNet-Parallel, MFISNet-Fused, and MFISNet-Refinement models tested have 3 1D convolutional layers followed by 3 2D convolutional layers; ReLU activations are used between layers. In the FYNet, MFISNet-Parallel and MFISNet-Refinement models, we use 1D and 2D convolutional kernels with $24$ channels following [Fan and Ying, 2022]; in the MFISNet-Fused models, we use 1D and 2D convolutional kernels with $24N_{k}$ channels to adjust for the increased input size. To train MFISNet-Fused, MFISNet-Parallel and MFISNet-Refinement, we search over a grid of architecture and optimization hyperparameters. We report the optimal hyperparameters we found in Tables 6, 9, 8, 7 and 11.
The FYNet model and all MFISNet models use no input or output normalization. The parameters for the 1D convolutional layers were initialized by drawing from a uniform distribution $U[0,2/d_{in}]$, where $d_{in}$ is the input dimension for a given layer. The parameters for the 2D convolutional layers were initialized by the standard initialiaztion scheme for 2D convolutional layers in Pytorch, We performed preliminary investigation into different initialization schemes and found that the initialization scheme had a smaller effect than architecture or optimization hyperparameters. As a result, we decided to restrict our hyperparameter search to the architecture and optimization hyperparameters.
@@ B.1 FYNet and MFISNet Models
1d kernel size 60 60
2d kernel size 5 5
Weight decay $0.0$ $0.0$
@@
1d kernel size $40$ $20$ $40$ $40$ $40$ $20$ $40$ $20$
2d kernel size $7$ $5$ $5$ $5$ $5$ $5$ $5$ $5$
Weight decay $1\times 10^{-3}$ $0.0$ $1\times 10^{-3}$ $0.0$ $0.0$ $0.0$ $1\times 10^{-3}$ $1\times 10^{-3}$
Learning rate $5\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
LR decrease $0.25$ $1.0$ $0.25$ $1.0$ $0.25$ $1.0$ $1.0$ $1.0$
@@
1d kernel size $40$ $20$ $20$ $20$ $40$ $40$ $20$ $20$
2d kernel size $5$ $7$ $5$ $7$ $5$ $5$ $7$ $7$
Weight decay $0.0$ $1\times 10^{-3}$ $0.0$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $0.0$
Learning rate $5\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$
@@
1d kernel size 40 20 20 40 20 20 20 40
2d kernel size 5 5 5 5 7 5 5 5
Weight decay 0 $1\times 10^{-3}$ 0 0 $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Learning Rate $5\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-3}$
@@
Hyperparameter Data Setting (MFISNet-Refinement)
with No Progressive Refinement Training
$N_{k}=2$ $N_{k}=3$ $N_{k}=4$ $N_{k}=5$
1d kernel size $40$ $20$ $40$ $40$
2d kernel size $5$ $5$ $5$ $7$
Weight decay $0.0$ $1\times 10^{-3}$ $1\times 10^{-3}$ $0.0$
Learning rate $5\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Table 10: Optimal hyperparameters for the MFISNet-Refinement models trained with the “No Progressive Refinement” training condition (Section 5.5). All of the models were trained on noiseless training samples.
Hyperparameter Data Setting (MFISNet-Refinement)
with No Homotopy through Frequency Training
$N_{k}=2$ $N_{k}=3$ $N_{k}=4$ $N_{k}=5$
1d kernel size $40$ $20$ $40$ $40$
2d kernel size $5$ $7$ $7$ $7$
Weight decay $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Learning rate $5\times 10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-4}$
$\mathbf{\gamma}$ $1.0$ $1.0$ $1.1$ $1.1$
Table 11: Optimal hyperparameters for the MFISNet-Refinement models trained with the “No Homotopy through Frequency” training condition (Section 5.5). Here, $\mathbf{\gamma}$ is the factor which weights different loss terms (cf. (18)). We searched over values $\gamma=\{0.9,1.0,1.1\}$. All of the models were trained on noiseless training samples.
@@ B.2 Wide-Band Butterfly Network
Appendix C Training Method with Warm-Start Initialization
In this section, we give details of the “Algorithm 2 + Warm-Start Initialization” training procedure described in Section 5.6. This training procedure leverages the intuition that sequential blocks in our MFISNet-Refinement architecture learn similar functions. In Algorithm 3 we give full pseudocode for the warm-start training procedure. In this pseudocode, we refer to the trainable parameters in block $t$ as $\theta_{t}$. If $t>1$, these parameters contain the parameters for the FYNet block, which we call $\theta_{t,1}$ and the parameters for the extra filtering layers, which we call $\theta_{t,2}$.
Input: Training data samples $\mathcal{D}_{n}:=\left\{(q^{(j)},d_{k_{1}}^{(j)},\ldots,d_{k_{N_{k}}}^{(j)})\right\}_{j=1}^{n}$.
1 for $t=1,...,N_{k}$ do
2      if $t=1$ then
3           Initialize $t_{1,1}$ randomly
4           else if $t=2$ then
5                Initialize $\theta_{2,1}$ with $\theta_{1,1}$
6                Initialize $\theta_{2,2}$ randomly
7                else
8                     Initialize $\theta_{t,1}$ with $\theta_{t-1,1}$
9                     Initialize $\theta_{t,2}$ with $\theta_{t-1,2}$
10
11                     Set $\theta_{t}$ as trainable, and freeze all other weights
12                     if $t<N_{k}$ then
                          Train $\theta_{t}$ by optimizing $L_{t}$
                           // Equation 14
13
14                          else
15                               Train $\theta_{t}$ by optimizing $\|\hat{q}_{k_{N_{k}}}-q\|_{2}^{2}$
16
17                               Set all weights as trainable
18                               Train all weights by optimizing $\|\hat{q}_{k_{N_{k}}}-q\|_{2}^{2}$
Algorithm 3 Training Procedure
Appendix D Additional empirical results
In this section, we illustrate the predictions generated by the different models used in our experiments on randomly-selected test samples from our dataset. For each prediction, we include the associated error plot. In  Figure 9, we provide more samples comparing FYNet, Wide-Band Butterfly Network, and MFISNet-Refinement. In  Figure 10, we show sample outputs from MFISNet-Fused and MFISNet-Parallel.
Figure 9: Sample predictions from four randomly-selected test samples. The first row shows the ground-truth scattering potential; in this plot we show the wavelength corresponding to the maximum frequency $k=32\pi$.
Figure 10: Sample predictions from four randomly-selected test samples. The first row shows the ground-truth scattering potential; in this plot we show the wavelength corresponding to the maximum frequency $k=32\pi$.